# CAE-Net: Generalized Deepfake Image Detection using Convolution and Attention Mechanisms with Spatial and Frequency Domain Features

### v2

@@ 3 Experimental Results
[…]
@@
Finally, we look into the effectiveness of our ensemble method through a simple adversarial perturbation attack, such as Fast Gradien Sign Method (FGSM). As mentioned before, FGSM can fool models by adding simple perturbation noise based on the gradient sign that intends to maximize the loss function of the detection model instead of minimizing it. We present the accuracy scores of the separate detection model as well as the ensemble models for different values of $\epsilon$ in Table 8.
Table 8: Comparison of Our Proposed Architectures under Adversarial FGSM Attacks
Model Accuracy
$\epsilon=0.01$ $\epsilon=0.03$ $\epsilon=0.05$
EfficientNetB0 38.54% 46.48% 53.94%
Data Efficient Image Transformer 42.76% 21.94% 16.73%
ConvNeXt with wavelet 6.45% 6.51% 15.40%
Weighted Ensem.(EfficientNet, DeiT, ConvNeXt) 12.04% 13.48% 16.05%
Majority Voting (EfficientNet, DeiT, ConvNeXt) 25.52% 20.41% 19.76%
@@
The results show an increased robustness of the EfficientNet architecture, which depicts increased accuracy with increasing perturbation compared to the other two architectures. DeiT shows a decrease in accuracy with rising perturbation, and ConvNeXt accuracy drops significantly even for very low perturbation added to the image. Though a weighted ensemble model is appropriate for achieving high accuracy on the generalized dataset, we argue that it is highly vulnerable to adversarial attacks. If any model performs poorly on adversarial data, then the performance of the entire ensemble drops significantly. However, in the majority voting ensemble, the performance does not reduce as much, since it requires the majority of models to perform poorly, implying increased robustness against adversarial attacks for applications, such as deepfake detection.
@@ 4 Conclusion
In this work, we proposed an ensemble architecture for deepfake detection that captured the intricate properties of the input images to successfully distinguish between real and fake images and produced excellent results. We emphasize that detecting deepfake images that are generated with a range of generative AI approaches is a challenging task. To that end, we have addressed a significant class imbalance through disjoint training set-based multistage training, and proposed a weighted ensemble approach that takes advantage of the strengths of an EfficientNet model, a Data Efficient Image Transformer model, and a ConvNext model for the accurate classification of deepfake images. The EfficientNet-B0 model enabled extracting local features from an image, while the DeiT model specialized in capturing global features. Additionally, ConvNeXt, combined with the wavelet transform, effectively identified frequency-dependent artifacts. Through t-SNE plots, we demonstrated that our model is capable of separating real and fake images into separable clusters. Additionally, we included Grad-CAM visualizations to map the focus regions of images to illustrate how our models analyze the images and which portions contribute to distinguishing between real and fake images. Besides, we further analyzed the effectiveness of the proposed models by testing them against FGSM attacks and found the optimum ensemble mechanism to be majority voting compared to the weighted ensemble with a little trade-off in accuracy. An interesting future work could be incorporating lightweight architectures, reducing the computational overhead, and increasing the model’s adaptability to newer and sophisticated deepfake generation methods.

### v3

@@ 3.5 Comparison with the existing works
  Detection Method Datasets AUC Score Std. Dev.
  Capsule [69] UADFV [34], DF-TIMIT [70], FaceForensics++ [31], 69.4% 13.84%
@@
We compared our proposed model with existing ensemble and frequency–spatial hybrid architectures on the DF-Wild Cup dataset to evaluate both classification performance and model complexity. As shown in Table 14, CAE-Net achieves superior performance, attaining 94.46% accuracy with 117.4M parameters, outperforming previous methods, such as ResNeXt–Xception, DeepfakeStack, and CoAt-Net ensembles. Unlike prior approaches that primarily focus on either spatial or hybrid CNN–Transformer integration without frequency-domain enhancement, CAE-Net incorporates wavelet-based preprocessing to extract complementary frequency cues while maintaining architectural diversity among its sub-models. This balanced design allows the ensemble to learn richer representations of deepfake artifacts and to generalize more effectively across diverse manipulation techniques. Consequently, CAE-Net achieves a better trade-off between accuracy and computational cost, demonstrating the effectiveness of its ensemble logic compared to existing state-of-the-art models.
Table 14: Comparative analysis of our work with prior ensembles and frequency-spatial hybrid networks
Classifiers # of params. Accuracy
Freq-Net [73] 1.85 M 67.35%
FSBI [21] 23.34 M 82.00%
CoAt-Net Ensemble [74] 699.72 M 84.22%
SFA-Net [75] 25.04 M 90.07%
ResNeXt & Xception Ensemble [3] 43.8 M 91.21%
DeepfakeStack [76] 135.55 M 92.12%
Proposed CAE-Net 117.4 M 94.46%
3.6 Robustness analysis against adversarial attacks
@@
Finally, we look into the effectiveness of our ensemble method through a simple adversarial perturbation attack, such as Fast Gradient Sign Method (FGSM). As mentioned before, FGSM can fool models by adding simple perturbation noise based on the gradient sign that intends to maximize the loss function of the detection model instead of minimizing it. We present the accuracy scores of the separate detection model as well as the ensemble models for different values of $\epsilon$ in Table 15.
Table 15: Comparison of our proposed architectures under FGSM attacks
Model Accuracy (before adversarial training) Accuracy (after adversarial training)
$\epsilon=0.005$ $\epsilon=0.01$ $\epsilon=0.03$ $\epsilon=0.05$ $\epsilon=0.00$ $\epsilon=0.005$ $\epsilon=0.01$ $\epsilon=0.03$ $\epsilon=0.05$
EfficientNetB0 42.77% 38.54% 46.48% 53.94% 93.07% 56.90% 50.39% 45.21% 50.26%
DeiT 49.61% 42.76% 21.94% 16.73% 89.29% 65.10% 37.14% 5.57% 3.74%
ConvNeXt with wavelet 16.08% 6.45% 6.51% 15.40% 92.25% 98.40% 96.45% 63.28% 51.63%
Weighted Ensem.(CAE-Net) 20.64% 12.04% 13.48% 16.05% 94.43% 87.27% 69.92% 14.84% 20.25%
Majority Voting (CAE-Net) 34.86% 25.52% 20.41% 19.76% 93.85% 82.58% 64.29% 25.03% 35.52%
@@
The results show an increased robustness of the EfficientNet architecture, which depicts increased accuracy with increasing perturbation compared to the other two architectures. DeiT shows a decrease in accuracy with rising perturbation, and ConvNeXt accuracy drops significantly even for very low perturbations added to the image, which is likely caused by the modification of frequency domain artifacts by even small attacks, which disrupts its feature extraction. Though a weighted ensemble model is appropriate for achieving high accuracy on the generalized dataset, we argue that it is highly vulnerable to adversarial attacks. If any model performs poorly on adversarial data, then the performance of the entire ensemble drops significantly. However, in the majority voting ensemble, the performance does not reduce as much, since it requires the majority of models to perform poorly, implying moderate robustness against adversarial attacks for applications, such as deepfake detection.
To improve the robustness of our models against adversarial attacks, we integrated adversarial training using pre-trained weights optimized for clean data. This involved training with adversarial samples generated using a small perturbation ($\epsilon=0.005$) over six epochs, employing a reduced learning rate of 1e-5 and a batch size of 64. We adopted a mixed batch strategy, comprising 32 perturbed images and 32 unperturbed images per batch, to balance exposure to both conditions [57]. Notably, this approach yielded significant enhancements, with models achieving improved accuracy on the clean validation set and demonstrating superior resilience to smaller FGSM perturbations ($\epsilon=0.005,0.01$) without adversarial training. The weighted ensemble and majority voting also show enhanced resilience to the FGSM attacks, with accuracy shifting from 20.64% without adversarial training to 87.27% (see Table 15) after adversarial training for $\epsilon=0.005$. The increase in overall accuracy is maintained for higher perturbations as well. Again, majority voting outperforms weighted ensemble in the case of higher perturbations, which is expected. Further training with incrementally larger perturbations and with more sophisticated adversarial images (such as, projected gradient descent) may enhance performance against higher perturbation levels, a direction we intend to explore in future work.
4 Discussion
This study presents CAE-Net, an ensemble architecture combining convolutional and attention-based models with spatial and frequency domain features for generalized deepfake detection. In this section, we discuss our design decisions, analyze the deployment implications of our results, acknowledge the limitations of our work, and outline future research directions.
The first major challenge of our study was addressing the 5:1 class imbalance in our dataset. Initially, we attempted conventional techniques like random oversampling and weighted loss functions. However, these techniques presented critical limitations, as stated in Section 2.2.1. Our method ensures balanced 1:1 exposure at each of the five training stages, with each stage using all 42,690 real images paired with a disjoint subset of 43,894 fake images. Through this method, the authenticity of real images is preserved, and knowledge accumulation is continued via warm starting the next stages. Recent estimates suggest AI-generated content now exceeds human-created content in volume online [77]. As deepfake generation becomes more accessible, detection datasets will increasingly exhibit class imbalance, with synthetic samples vastly outnumbering authentic ones. Our disjoint multistage training method provides a scalable solution for such scenarios without requiring proportional increases in authentic training data.
We position CAE-Net for deployment scenarios where accuracy is very important and computational resources are available, such as centralized content moderation systems or forensic laboratories. For resource-constrained environments, such as edge devices, the individual EfficientNet model provides a viable alternative with only 3.96 M parameters and 386 M FLOPs, as shown in Table 9. For scenarios where resource constraints are more relaxed, the ensemble of EfficientNet and ConvNeXt is also a good option with a 1.5s inference time and 31.77 M parameters with a high accuracy of 93.14%.
Note that the ensemble of EfficientNet and ConvNeXt provides detection accuracy comparable to CAE-Net with less than one-fourth the parameters and half the inference time. However, we emphasize here the 1.32% accuracy improvement of CAE-Net is equivalent to correctly identifying 13,200 additional deepfakes per million processed images in content moderation systems. Given the serious consequences of deepfake spread, including financial fraud, electoral interference, and non-consensual imagery, the requirement for high accuracy justifies our ensemble approach’s computational expense in critical applications.
Our evaluation under FGSM attacks revealed that CAE-Net’s accuracy drops significantly under perturbations for the weighted ensemble, while majority voting performs slightly better under the same attack. This vulnerability stems from our probability-level fusion mechanism: when adversarial perturbations cause even one model’s output probabilities to shift significantly, the weighted sum propagates this error proportionally. However, the binary decision mechanism of majority voting scheme provides a bounded impact where a single compromised model cannot override the other two. This represents a fundamental trade-off between clean-data optimization (weighted ensemble) and adversarial resilience (majority voting). However, upon adversarial training, the models show better robustness against such attacks, outlining the importance of such enhancement methods for deepfake classifiers. For deployments in adversarially exposed environments such as public-facing platforms, we recommend majority voting. For accuracy-critical applications with lower adversarial threat, the weighted ensemble remains preferable.
CAE-Net employs a probability-level fusion rather than intermediate feature-level fusion. Each model (EfficientNet, ConvNeXt, DeiT) processes inputs independently through its complete architecture, maintaining native feature representations without cross-model feature map alignment or shared layers. This design choice offers key advantages: architectural independence allows each model to be trained and optimized separately, computational efficiency is obtained through parallel inference, and the modularity of late fusion enables straightforward addition or removal of ensemble members. Fusion occurs only at the final layer through weighted averaging of class probabilities, making the ensemble mechanism interpretable and the contribution of each model directly quantifiable.
Limitations and Future Directions. Several limitations of our work require discussion. First, the dataset provided by the IEEE SPCup 2025 organizers had a fixed validation set. While defining the weights of the ensemble, we could not implement cross-validation due to this fixed split, which limits the robustness of our weight selection. Second, hyperparameter optimization was performed manually due to resource constraints rather than through systematic grid search or automated methods. Third, while testing the benchmark model performances described in Table 456, we used a fixed random seed of 42 for reproducibility and fair comparison. While this ensures consistent results for that specific initialization, it does not guarantee that our observed improvements generalize across different random initializations.
The Grad-CAM visualizations indicate that the models sometimes concentrate on irrelevant areas, such as the background or non-facial features, instead of the artifact-prone facial regions essential for accurate classification. This misdirected focus contributes to misclassifications. To address this issue, incorporation of attention regularization techniques can work to guide the models toward critical facial areas, prioritizing zones with significant artifacts. Additionally, our evaluation does not assess performance on compressed or degraded images, which are common in real-world social media contexts. The robustness of frequency-domain features to such compression artifacts requires further investigation. Our ablation studies focused on architectural combinations but did not explore alternative fusion strategies beyond weighted averaging and majority voting, such as learned meta-classifiers, attention-based fusion, or stacking. Future research directions include developing efficient model compression techniques, such as knowledge distillation, quantization, and pruning, to enable deployment on resource-constrained devices while maintaining high accuracy, and testing generalization to emerging deepfake generation methods, including diffusion models and neural radiance fields (NeRFs) [78].
@@ 5 Conclusion
In this work, we proposed an ensemble architecture for deepfake detection that captured the intricate properties of the input images to successfully distinguish between real and fake images and produced excellent results. We emphasize that detecting deepfake images that are generated with a range of generative AI approaches is a challenging task. To that end, we have addressed a significant class imbalance through disjoint training set-based multistage training and proposed a novel weighted ensemble approach that takes advantage of the strengths of an EfficientNet model, a Data Efficient Image Transformer model, and a ConvNext model for the accurate classification of deepfake images. The EfficientNet-B0 model enabled extracting local features from an image, while the DeiT model specialized in capturing global features. Additionally, ConvNeXt, combined with the wavelet transform, effectively identified frequency-dependent artifacts. Statistical analysis shows that the CAE-Net significantly outperforms the constituent backbone models. Through t-SNE plots, we demonstrated that our model is capable of separating real and fake images into separable clusters. Additionally, we included Grad-CAM visualizations to map the focus regions of images to illustrate how our models analyze the images and which portions contribute to distinguishing between real and fake images. Besides, we further analyzed the effectiveness of the proposed models by testing them against FGSM attacks and found the optimum ensemble mechanism to be majority voting compared to the weighted ensemble with a little trade-off in accuracy. An interesting future direction could be the incorporation of lightweight architectures, which would reduce computational overhead and increase the model’s adaptability to newer and more sophisticated deepfake generation methods.